Reduce the following fractions to the simplest form: 
84 / 98 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

84 / 98 = (14 × 6) / (14 × 7)
= 6/7

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The following fraction...
84/98  can be reduced as 
(84 ÷ 14)/ (98 ÷ 14)
= 6/7
 
Hope it helps!
Thank you